Please, Calculate the equation of the linear function which contains the ordered pairs (-4,1) and (3,6).
Answer by checkley71(8403)

Y=MX+B OR Y=(1-6)/(3-(-4))+B OR Y=(-5/1)X+B OR Y=-5X+B